# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ Abstract
Information criteria (IC) have been widely used in factor models to estimate an unknown number of latent factors. It has recently been shown that IC perform well in Common Correlated Effects (CCE) and related setups in selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors. As CAs can proxy non-stationary factors, it is tempting to claim such generality of IC, too. We show formally and in simulations that IC have a severe underselection issue even under very mild forms of factor non-stationarity, which goes against the sentiment in the literature.
@@ 1 Introduction
Consider the following set of $K$ variables that admits a factor structure for $i=1,\ldots,N$:
@@
where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is the error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ is latent common factors that can be either stationary or not, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix. (1.1) nests several popular settings.
@@
Example 1. Common Correlated Effects (CCE). Let $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables and $\mathbf{y}_{i}$ is the dependent variable. Here, $K=k+1$, and $i$ indexes individuals, such as countries or households. Then (1.1) represents a classical CCE setup of Pesaran, (2006) with $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ and $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$. Here, $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter vector, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings). Both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, and this induces cross-section dependence among individuals and endogeneity as the error in $\mathbf{y}_{i}$ is correlated with $\mathbf{X}_{i}$. (1.1) nests this model with $\mathbf{C}_{i}=[\boldsymbol{\Gamma}_{i}\boldsymbol{\beta}+\boldsymbol{\gamma}_{i},\boldsymbol{\Gamma}_{i}]$ and $\mathbf{U}_{i}=[\mathbf{V}_{i}\boldsymbol{\beta}+\boldsymbol{\varepsilon}_{i},\mathbf{V}_{i}]$.
Example 2. Distinct Correlated Effects (DCE). We can introduce distinct but correlated factors $\mathbf{F}=[\mathbf{F}_{\mathbf{x}},\mathbf{F}_{\mathbf{y}}]\in\mathbb{R}^{T\times(m_{x}+m_{y})}$ in Example 1, which makes the setup mode applicable: $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}_{\mathbf{y}}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$, $\mathbf{X}_{i}=\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$ as considered in Cui et al., (2022) or Stauskas and De Vos, (2025). Here, (1.1) nests such model similarly to Example 1, but $\boldsymbol{\gamma}_{i}$ is modified accordingly, depending on whether $\mathbf{F}_{\mathbf{y}}\subset\mathbf{F}_{\mathbf{x}}$, vice versa or $\mathbf{F}_{\mathbf{x}}\cap\mathbf{F}_{\mathbf{y}}=\emptyset$.
Example 3. Block Structure (FRED-MD-QD , EA-MD-QD). The regressors in Examples 1 and 2 can be high-dimensional, but follow a natural structure of $k$ blocks and $m$ underlying factors. Let $i_{j}$ index series in block $j$ for $j=1,\ldots,k$. Then $\mathbf{Z}_{i}=\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}=[\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{k}}]^{\prime}$. Here, $N$ is the number of series per block. Monetary, industrial or fiscal blocks of series can be found in FRED-MD (see McCracken and Ng,, 2016), FRED-QD (see McCracken and Ng,, 2020) or EA-MD-QD (see Barigozzi et al.,, 2024) datasets. These datasets have been very popular in factor-augmented forecasts (see e.g. Karabiyik and Westerlund,, 2021, or Morico and Stauskas,, 2025).
In all the cases above, we can estimate the factors by taking $K$ cross-section averages (CAs): $\widehat{\mathbf{F}}=\overline{\mathbf{Z}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g. Pesaran and Tosetti,, 2011). In Example 1, $\widehat{\mathbf{F}}$ is used as an additional regressor to merely control for the factors, and subsequently $\boldsymbol{\beta}$ is estimated with the least squares (LS), which constitutes the typical CCE estimator (see e.g. Karabiyik et al.,, 2017). In Example 2, only $\widehat{\mathbf{F}}_{\mathbf{x}}=\overline{\mathbf{X}}$ is used as a regressor, while $\mathbf{F}_{\mathbf{y}}$ is controlled by bootstrap (see Stauskas and De Vos,, 2025). In Example 3, the coefficient of $\widehat{\mathbf{F}}$ itself matters for forecast evaluation (see e.g. Massacci and Kapetanios,, 2024).
@@
A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties. Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings without any modifications (see e.g. Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023). This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs ($K\in\{k,k+1\}$ in Examples 1-3). However, there are reasons to ensure that $m=K$. If $m<K$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017). This situation is common in macro datasets (see De Vos and Stauskas,, 2024). In DCE, $\mathbf{F}_{\mathbf{x}}$ estimated with an excess error makes bootstrap inconsistent (see Stauskas and De Vos,, 2025). In Example 3, popular tests of forecast accuracy comparison have highly non-standard asymptotic distributions (see Stauskas and Westerlund,, 2022). In general, too many CAs might even lead to a model that over-fits the data. In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion). In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that Information Criteria (IC) of Bai and Ng, (2002) from Principal Components (PC) literature can be applied in the CCE setting in Example 1 under stationary factors to identify an optimal set of CAs. De Vos and Stauskas, (2024) (DVS) showed similar results applicable to more general Examples 2 and 3.111Juodis, (2022) considers eigenvalue-based methods of Ahn and Horenstein, (2013)) under stationarity. The focus there is the number of factors but not the optimal set of CAs. We leave analysis of this approach for the future research.
MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC. The power of CAs to proxy general factor structures makes it natural to evaluate this statement and re-visit IC proposed in MW and DVS. We show that, contrary to the sentiment in the CCE/CAs literature, IC crumble under very mild forms of nostationarity. We demonstrate formally and in simulations that they exhibit a severe under-selection problem, which is independent of penalties used in the objective functions of IC. Moreover, nonstationarity of $\mathbf{F}$ cannot be accommodated in the objective functions a priori, as their very design is based on the idea that properties of factors and idiosyncratics should be similar. This study can be seen as a caution to practitioners to transform data and not to extrapolate attractive properties of CAs should they need to choose a number thereof.
@@ 2 Econometric Setup
To further operationalize our examples, we introduce $M$ which is a set of column indices of $\overline{\mathbf{Z}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{K\times g}$ picks the corresponding $g$ averages in practice. For instance, $\overline{\mathbf{Z}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k+1$ CAs in the typical CCE setup. Moreover, let $\mathbf{q}_{M}=[\boldsymbol{0}_{g\times 1},\mathbf{q}_{M_{\mathbf{x}}}^{\prime}]^{\prime}$, then $\overline{\mathbf{Z}}\mathbf{q}_{M}=\overline{\mathbf{X}}\mathbf{q}_{M_{\mathbf{x}}}=\widehat{\mathbf{F}}_{M_{\mathbf{x}}}$ selects only from the set of $k$ regressors (or blocks), as required in Examples 2 and 3.222Throughout the remainder of the study, we use $m$ to denote the number of factors without the loss of generality. Clearly, this number can vary among our examples. Let accordingly $M_{0}$ ($M_{\mathbf{x},0}$) denote the set of averages from $\overline{\mathbf{Z}}$ ($\overline{\mathbf{X}}$) such that $\mathrm{rank}(\overline{\mathbf{C}}\mathbf{q}_{M_{0}})=\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x},0}})=m$, and $|M_{0}|=|M_{\mathbf{x},0}|=m$, where $|M|$ denotes cardinality of an arbitrary set $M$. IC under consideration are given by
$\mathrm{IC}^{MW}(M)=\ln\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}\right)+g\cdot p_{N,T},$ (2.1)
$\mathrm{IC}^{DVS}(M_{\mathbf{x}})=\ln\det\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}\mathbf{X}_{i}\right)+g\cdot k\cdot p_{N,T},$ (2.2)
$p_{N,T}\in\left\{\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right)\right\}$ (2.3)
where $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix onto the orthogonal complement of $\mathbf{A}$, $\mathbf{A}^{+}$ is a Moore-Penrose (MP) inverse, while $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ with $\widehat{\boldsymbol{\beta}}$ is obtained with the CCE estimator under all available $k+1$ CAs. Finally, $p_{N,T}$ is a penalty term which goes to 0 with $(N,T)\to\infty$. Examples of feasible and most popular penalties are given in (2.3) for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002). Note that (2.1) - (2.2) are functions of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
Let $\overline{M}$ ($\overline{M}_{\mathbf{x}}$) be the set of indices of all available CAs that $|\overline{M}|=k+1$ ($|\overline{M}_{\mathbf{x}}|=k$). Then
$\widehat{M}\hskip 2.84526pt(\widehat{M}_{\mathbf{x}})=\operatorname*{arg\,min}_{M\subseteq\overline{M}\hskip 2.84526pt(M_{\mathbf{x}}\subseteq\overline{M}_{\mathbf{x}})}\mathrm{IC}(M)\hskip 2.84526pt(\mathrm{IC}(M_{\mathbf{x}})),$ (2.4)

### v3

@@ Abstract
Information criteria (ICs) have been widely used in factor models to estimate an unknown number of latent factors. It has recently been shown that ICs perform well in Common Correlated Effects (CCE) and related settings when selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors. As CAs can proxy non-stationary factors, it is tempting to claim an excellent performance of ICs under general factors, too. We show formally and in simulations that they remain consistent, but the more persistent factors are, the poorer they perform in small samples, which goes against the sentiment in the CCE literature.
Abstract
In this Supplementary Material, we provide the proofs of our auxiliary results and our main result in Proposition 1 (Sections 2 and 3). In addition, we demonstrate how the proofs change in the case of the IC of Margaritella and Westerlund, (2023). In Section 4, we provide additional simulation evidence, including, e.g., correlated idiosyncratics or eigenvalue ratio method of Juodis, (2022).
@@ 1 Introduction
Consider a set of $K$ variables (stacked over time $t=1,\ldots,T$) that admits a factor structure for cross-sectional units $i=1,\ldots,N$:
@@
where $\mathbf{Z}_{i}=[\mathbf{z}_{i,1},\ldots,\mathbf{z}_{i,T}]^{\prime}\in\mathbb{R}^{T\times K}$, $\mathbf{U}_{i}$ is an error term, while $\mathbf{F}\in\mathbb{R}^{T\times m}$ are latent common factors, where $m$ denotes the number of factors, and $\mathbf{C}_{i}\in\mathbb{R}^{m\times K}$ is the loading matrix. The most popular setting nested in (1.1) is that of interactive effects. Then $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, and $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ gives a model with unobserved heterogeneity, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables. Naturally, $K=k+1$, and $i$ indexes individuals, such as countries or households. The unobserved component of $\mathbf{y}_{i}$ is likely to be correlated with $\mathbf{X}_{i}$ (see Bai, 2009a, ). If $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, (1.1) represents the Common Correlated Effects (CCE) setting of Pesaran, (2006), where $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter of interest, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings). Since both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, we have endogeneity and cross-section dependence among individuals that make the least squares (LS) estimator of $\boldsymbol{\beta}$ inconsistent.
It is possible to estimate unobserved factors in several ways, including the Principal Components (PC) method (see Bai and Ng,, 2006) or diversified projections (see Fan and Liao,, 2022). Pesaran, (2006) suggests a simple and elegant way to proxy $\mathbf{F}$ up to a linear transformation by taking cross-section averages (CAs): $\overline{\mathbf{Z}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]=\widehat{\mathbf{F}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g. Pesaran and Tosetti,, 2011). The CCE estimator is then given by
$\widehat{\boldsymbol{\beta}}_{CCE}=\overline{\mathbf{Q}}^{-1}\overline{\mathbf{q}},$ (1.2)
where $\overline{\mathbf{Q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{X}_{i}$, $\overline{\mathbf{q}}=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}}\mathbf{y}_{i}$, $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix and $\mathbf{A}^{+}$ is the Moore-Penrose inverse. The core principle of CCE is to project the variables onto the orthogonal complement of their CAs in order to purge the factors asymptotically. Note that CAs are applicable beyond the usual interactive effects and CCE. Stauskas and De Vos, (2025) consider Distinct Correlated Effects (DCE), where $\mathbf{y}_{i}$ and $\mathbf{X}_{i}$ load on distinct sets of factors, and (1.2) is aided by bootstrap. Massacci and Kapetanios, (2024) use $\overline{\mathbf{X}}$ to approximate factors and produce factor-augmented forecasts.
@@
A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties. Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings without any modifications (see e.g. Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023). This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs. However, there are reasons to ensure that $m=k+1$. If $m<k+1$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017). This situation is common in macro datasets (see De Vos and Stauskas,, 2024). In general, too many CAs might even lead to a model that over-fits the data. In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion). Beyond CCE, Stauskas and De Vos, (2025) show that DCE only uses $\overline{\mathbf{X}}$, but $m=k$ must hold. In forecasting, the same is needed to avoid only conservative confidence intervals around factor-augmented forecasts (see Karabiyik and Westerlund,, 2021). In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that the information criteria (ICs) of Bai and Ng, (2002) and Bai, 2009b from the PC literature can be applied in the CCE setting under stationary factors to identify an optimal set of CAs from $\overline{\mathbf{Z}}$. De Vos and Stauskas, (2024) (DVS) show similar results when CAs are selected from $\overline{\mathbf{X}}$, which is more applicable in the DCE and forecasting settings. [^1] Even in basic CCE setting, it is advised to omit $\overline{\mathbf{y}}$ and use $\overline{\mathbf{X}}$ only to avoid computational issues (see Karavias et al.,, 2023)., [^2] PC techniques are limited in CCE, as they focus solely on $m$ by detecting the largest eigenvalues of the data matrix. They cannot detect sets of CAs as they do not have a natural ordering. With IC, we learn $m$ from the cardinality of the selected set.
MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC. As CAs proxy a general factor structure, it is natural to evaluate this statement and re-visit the ICs proposed in MW and DVS. Therefore, this study can be seen as the CAs counterpart of Bai, (2004), where classical PC results of Bai and Ng, (2002) are evaluated against pure unit root factors. Instead, we use a mildly integrated process by Magdalinos and Phillips, (2009) to model $\mathbf{F}$, which allows us to experiment with varying degrees of factor persistence. We demonstrate formally and in simulations that, while ICs remain consistent as $(N,T)\to\infty$, highly persistent factors negatively affect their small sample performance. We discuss differences between our results and those in Bai, (2004), and also explore penalty functions adapted to non-stationary $\mathbf{F}$, as suggested in the latter study. Since they render our ICs inconsistent, we argue that practitioners should not mechanically take recommendations from the PC literature if the goal is to select CAs. More importantly, the ICs of DVS and MW (with the original penalties) should be applied in relatively large samples if the presence of non-stationary factors is suspected.
@@ 2 Econometric Setup
We focus on the IC of DVS due to its wider applicability. While we provide numerical experiments on both DVS and MW, theoretical arguments regarding the latter are nearly identical, and we relegate them to Section 3.3 of the Supplementary material. To operationalize our analysis, we introduce $M$ which is a set of column indices of $\overline{\mathbf{X}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{k\times g}$ picks the corresponding $g$ averages in practice. For instance, $\overline{\mathbf{X}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k$ CAs. Consequently, let $M_{0}$ denote the true set of averages from $\overline{\mathbf{X}}$ such that $\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}})=m$, and $|M_{0}|=m$, where $|M|$ denotes the cardinality of an arbitrary set $M$. IC under consideration is given by
$\mathrm{IC}(M)=\ln\det\left(\overline{\mathbf{Q}}_{M}\right)+g\cdot k\cdot p_{N,T},$ (2.1)
where $\ln(.)$ denotes the natural logarithm and $\mathrm{det}(\mathbf{A})$ is a determinant of any square matrix $\mathbf{A}$. $\overline{\mathbf{Q}}_{M}$ is $\overline{\mathbf{Q}}$ from (1.2) evaluated at $\widehat{\mathbf{F}}_{M}$, and $p_{N,T}$ is a penalty term. [^3] In case of MW, $\overline{\mathbf{Q}}_{M}=\ln\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M}}\widehat{\boldsymbol{\nu}}_{i}\right)$, where $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ and $\widehat{\boldsymbol{\beta}}$ is obtained using the CCE estimator under all available $k+1$ CAs. Examples of feasible and most popular penalties are given by
$p_{N,T,1}=\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\hskip 5.69054ptp_{N,T,2}=\frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right)$ (2.2)
for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002). Note that (2.1) is function of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
Let $\overline{M}$ be the set of indices of all available CAs that, so that $|\overline{M}|=k$. Then
$\widehat{M}\hskip 2.84526pt=\operatorname*{arg\,min}_{M\subseteq\overline{M}\hskip 2.84526pt}\mathrm{IC}(M),$ (2.3)